### v1

@@ Abstract
Category imbalance is one of the most popular and important issues in the domain of classification. In this paper, we present a new generalized framework with Adaptive Weight function for soft-margin Weighted SVM (AW-WSVM), which aims to enhance the issue of imbalance and outlier sensitivity in standard support vector machine (SVM) for classifying two-class data. The weight coefficient is introduced into the unconstrained soft-margin support vector machines, and the sample weights are updated before each training. The Adaptive Weight function (AW function) is constructed from the distance between the samples and the decision hyperplane, assigning different weights to each sample. A weight update method is proposed, taking into account the proximity of the support vectors to the decision hyperplane. Before training, the weights of the corresponding samples are initialized according to different categories. Subsequently, the samples close to the decision hyperplane are identified and assigned more weights. At the same time, lower weights are assigned to samples that are far from the decision hyperplane. Furthermore, we also put forward an effective way to eliminate noise. To evaluate the strength of the proposed generalized framework, we conducted experiments on standard datasets and emotion classification datasets with different imbalanced ratios (IR). The experimental results prove that the proposed generalized framework outperforms in terms of accuracy, recall metrics and G-mean, validating the effectiveness of the weighted strategy provided in this paper in enhancing support vector machines.
@@
[…]
@@ 1 Introduction
Our primary focus is on the fundamental problem of standard SVMs. Descent techniques can be employed to minimize the objective function of a conventional SVM, which is known for its high convexity. Nevertheless, descent methods necessitate accurate calculation of the gradient of the goal function, a task that is frequently arduous. The Stochastic Gradient Descent (SGD) method addresses this issue by employing unbiased gradient estimates using tiny data samples [17, 18]. It is the primary approach for solving large-scale stochastic optimization issues. Nevertheless, the practical attractiveness of SGD approaches remains restricted due to their need for a substantial number of iterations to achieve convergence. Indeed, stochastic gradient descent (SGD) exhibits sluggish convergence due to its reliance on gradients, a drawback that is further amplified when gradients are substituted with random approximations. Various stochastic optimization strategies have been employed thus far to address SVM models. Aside from first-order techniques such as stochastic gradient descent (SGD), second-order approaches are also extensively employed. Second-order stochastic optimization methods impose greater demands on functions compared to first-order methods, as they are capable of quadratic differentiation. Consequently, studying these approaches is a demanding task. Second-order stochastic optimization methods inherently address the issues of limited precision and sluggish convergence that are associated with first-order approaches. This paper considers solving SVM using second-order stochastic optimization methods to further reduce the impact of outliers and dataset imbalance while ensuring accuracy.
@@
In this paper, we introduce a novel generalized framework with adaptive weight function for imbalanced SVM classification (AW-WSVM). The approach involves assigning weights to distinct samples and updating these weights based on the position of the hyperplane. Our main contributions are as follows:
@@
1. This work presents a novel weighting technique for support vector machines. The contribution of a training sample to the optimal classification hyperplane is determined by its distance from the hyperplane. Specifically, the closer the sample is to the hyperplane, the higher its weight is set, whereas the weight is smaller for samples that are farther away. We suggest the implementation of a novel weight function. The weight function is adaptive, and it can adjust itself with the change of hyperplane. It is deemed appropriate to assume that the weight setting follows a Gaussian distribution and that the impact of distance on the weight is stressed.
2. We enhance the unconstrained soft margin support vector machine. When the weight increases, the hinge loss function approaches zero to minimize the objective function, thus dividing the samples correctly.
3. We preprocess the data in each iteration. According to the directionality of the distance, we can select the noise data according to certain rules and eliminate it in the next training.
@@
4. Employ the weighting technique we suggested in conjunction with various optimization algorithms, and conduct experiments using datasets with different characteristics. In the field of emotion classification, we have done many experiments under different imbalanced ratios. The results demonstrate that the suggested weighting technique can enhance the performance of SVM even more.
@@
The remaining sections of this paper are structured in the following manner. Section 2 provides a concise overview of the recent literature on the classification of imbalanced data. Section 3 presents the introduction of a novel weighted method. Section 4 involves conducting experiments on various datasets by integrating different algorithms with our suggested weighting technique. The discussion and conclusions can be found in section 5.
@@ 2.1 Imbalanced Data
Data imbalances occur in the real world, often leading to overfitting. Faced with the problem of imbalance in SVM, there are two popular solutions. The first is to reconstruct the training points. Increase the number of points of a few classes in different ways (oversampling) or decrease the number of points of a majority class (undersampling) [19, 20]. The second is to introduce appropriate weight coefficients into the loss function. The error cost or the decision threshold of imbalanced data is adjusted to increase the weight of the minority class and reduce the weight of the majority class to achieve balance. An initial approach to incorporating weighting into SVM is through the use of fuzzy support vector machine (FSVM) [21]. Fuzzy SVM assigns different fuzzy membership degrees to all samples in the training set, and retrains the classifier. Shao et al.[22] introduced an extremely efficient weighted Lagrangian twin support vector machine (WLTSVM) that utilizes distinct training points to create two nearby hyperplanes for the classification of unbalanced data. Fan et al. [23] proposed an improved method based on FSVM to determine fuzzy membership by class determinism of samples, so that distinct samples can provide different contributions to the classification hyperplane.Kang et al. [24] introduced a downsampling method for Support Vector Machines (SVM) that uses spatial geometric distances. This method assigns weights to the samples depending on their Euclidean distance from the hyperplane. The weights highlight the samples’ influence on the decision hyperplane.
@@ 3.1 Support Vector Machine(SVM)
$\mathop{\min}_{\mathbf{w}\in\mathbb{R}^{n}}F(\mathbf{w})$ $=\cfrac{1}{l}\sum\limits_{i=1}^{l}f(\mathbf{w},\theta)$ (3)
$=\cfrac{C}{2}\lVert\mathbf{w}\rVert^{2}+\cfrac{1}{l}\sum\limits_{i=1}^{l}l((x_{i},y_{i});\mathbf{w})$
where $f(\mathbf{w},\theta_{i})=\cfrac{C}{2}\lVert\mathbf{w}\rVert^{2}+\text{max}(0,1-y_{i}(\mathbf{w^{T}}x_{i}))$, and $C>0$ is a hyperparameter.
@@ 3.2 Soft-Margin Weighted Support Vector Machine (WSVM)
$\mathop{\min}_{\mathbf{w}\in\mathbb{R}^{n}}F(\mathbf{w})$ $=\dfrac{1}{l}\sum\limits_{i=1}^{l}f(\mathbf{w},\theta)$ (4)
$=\dfrac{C}{2l}\sum\limits_{i=1}^{l}\lVert\mathbf{w}\rVert^{2}+\dfrac{1}{l}\sum\limits_{i=1}^{l}l((x_{i},y_{i},\alpha_{i});\mathbf{w})$
where $f(\mathbf{w},\theta_{i})=\frac{1}{2}\alpha_{i}C\lVert\mathbf{w}\rVert^{2}+\text{max}(0,1-y_{i}(\mathbf{w^{T}}x_{i}))$ We need to ensure that the objective function $F(\mathbf{w})$ is minimized when $\alpha_{i}$ is larger. Since the second term of the objective function is zero, it can only be that the $l_{2}$ norm of $\mathbf{w}$ tends to be smaller. It satisfies the spacing maximization principle.
@@ 3.3 Adaptive Weight Function(AW Function)
The initial decision hyperplane $\mathbf{w_{0}}x+b=0$ is obtained. Set $d_{i}^{k}$ is the distance between the $i$ th sample and the initial decision hyperplane, in the $k$ th iteration, expression is as follows:
$d_{i}^{k}=\frac{|\mathbf{w_{k}}x_{i}+b|}{\lVert\mathbf{w_{k}}\rVert},i=1,…,l$ (6)
@@
$\alpha_{ki}=\frac{2}{\sqrt{2\pi}\sigma}e^{-\frac{(d_{i}^{k})^{2}}{2\sigma^{2}}}+\frac{1}{M-m}e^{-\frac{d_{i}^{k}}{M-m}}$ (7)
@@
Figure 4: AW function with different $\sigma$ ($\sigma\in\{0.5,1,1.5\}$)
@@
((a)) Weights of Different Samples
((b)) 3D Visualisation of Weights
Figure 5: Weights generated by AW function
@@ 3.4 Generalized Framework for AW-WSVM
All samples are initially assigned the same weight, i.e., $\alpha_{0i}=\frac{2}{l},i=1,...,l$, then $\sum\limits_{i=1}^{l}\alpha_{0i}=2$. The assumption in Section 3.3 is given to illustrate the rationality of the above statement.
Now, we can utilize the AW function mentioned before to produce weights for training the AW-WSVM algorithm. Enter the distance $d_{i}^{k}(i=1,2,...,l)$ obtained by the $kth$ training into the (7). Then update the weight of each sample $\alpha_{ki},i=1,...,l$. And then we put it in the objective function again. The framework includes the following steps:
@@
(2) Screen samples
We can calculate the distance (vector) between each sample and the hyperplane. Remember that the positive class sample set is $S_{+}$ and the negative class is $S_{-}$. If there is $x_{i}\in S_{+}$, so that the inner product of $x_{i}$ and $x_{j}$ (any $x_{j}\in S_{+}$) is less than or equal to 0, it means that $x_{i}$ is noise. We will eliminate it. In the same way, $S_{-}$ is also treated in this way;
@@
1:The training set $(x_{i},y_{i},\alpha_{ki}),i=1,…,l$, the initial weight $\alpha_{0}=[\frac{2}{l},…,\frac{2}{l}]^{T}$, the original $\mathbf{w_{0}}=[0,…,0]^{T}$ and $k_{max}$.
@@
6:     Calculate:
7:     $d_{i}^{k}=\frac{|\mathbf{w_{k}}x_{i}+b|}{\lVert\mathbf{w_{k}}\rVert},i=1,…,l$
8:     Eliminate noise
9:     set
@@
10:     Update $\alpha_{k}=[\alpha_{k1},…,\alpha_{kl}]^{T}$
$\alpha_{ki}=\frac{2}{\sqrt{2\pi}\sigma}e^{-\frac{(d_{i}^{k})^{2}}{2\sigma^{2}}}+\frac{1}{M-m}e^{-\frac{d_{i}^{k}}{M-m}}$
11:     $k\leftarrow k+1$
12:end while

### v3

@@ Abstract
Category imbalance is one of the most popular and important issues in the domain of classification. Emotion classification model trained on imbalanced datasets easily leads to unreliable prediction. The traditional machine learning method tends to favor the majority class, which leads to the lack of minority class information in the model. Moreover, most existing models will produce abnormal sensitivity issues or performance degradation. We propose a robust learning algorithm based on adaptive cost-sensitivity and recursive denoising, which is a generalized framework and can be incorporated into most stochastic optimization algorithms. The proposed method uses the dynamic kernel distance optimization model between the sample and the decision boundary, which makes full use of the sample’s prior information. In addition, we also put forward an effective method to filter noise, the main idea of which is to judge the noise by finding the nearest neighbors of the minority class. In order to evaluate the strength of the proposed method, we not only carry out experiments on standard datasets but also apply it to emotional classification problems with different imbalance rates (IR). Experimental results show that the proposed general framework is superior to traditional methods in Accuracy, G-mean, Recall and F1-score.
@@
Imbalanced data · Robust learner · Adaptive Cost-Sensitive Learning · Prior information · Recursive Denoising
@@ 1 Introduction
Our primary focus is on the fundamental problem of soft-margin SVM. Wang et al.[17] improved the soft-margin support vector machine and proposed a new fast support vector machine with truncated square hinge loss, which fully improved the calculation speed of large-scale datasets. Descent techniques can be employed to minimize the objective function of a conventional SVM, which is known for its high convexity. Nevertheless, descent methods necessitate accurate calculation of the gradient of the goal function, a task that is frequently arduous. The Stochastic Gradient Descent (SGD) method addresses this issue by employing unbiased gradient estimates using tiny data samples [18, 19]. It is the primary approach for solving large-scale stochastic optimization issues. Nevertheless, the practical attractiveness of SGD approaches remains restricted due to their need for a substantial number of iterations to achieve convergence. Indeed, SGD exhibits sluggish convergence due to its reliance on gradients, a drawback that is further amplified when gradients are substituted with random approximations. Various stochastic optimization strategies have been employed thus far to address SVM models. Aside from first-order techniques such as SGD, second-order approaches are also extensively employed. Second-order stochastic optimization methods impose greater demands on functions compared to first-order methods, as they are capable of quadratic differentiation. Consequently, studying these approaches is a demanding task. Second-order stochastic optimization methods inherently address the issues of limited precision and sluggish convergence that are associated with first-order approaches. This paper considers solving SVM using second-order stochastic optimization methods to further reduce the impact of outliers and dataset imbalance while ensuring accuracy.
@@
In this paper, we introduce a novel generalized framework with adaptive weight function for imbalanced SVM classification (AW-WSVM). Our main contributions are as follows:
@@
1. This work proposes a novel cost-sensitive learning technique. In the kernel space, the contribution of training samples to the optimal classification hyperplane is determined by their kernel distance from the hyperplane. Specifically, the closer a sample is to the hyperplane, the greater the penalty term for its loss, while the farther a sample is from the hyperplane, the smaller the penalty term for its loss. We suggest implementing a new adaptive weight function that can adjust itself according to the changes in the decision hyperplane.
2. We introduce a new soft-margin weighted support vector machine. The model no longer focuses on the overall loss but instead emphasizes the loss of each sample.
3. We filter the data during each iteration, paying more attention to the samples in the vicinity of the decision hyperplane. As a small number of samples are sensitive to noise, we propose a recursive method to find the nearest neighbors and eliminate noisy data.
@@
4. Employ the adaptive cost-sensitive learning and recursive denoising technique we suggested in conjunction with various optimization algorithms, and conduct experiments using datasets with different characteristics. In the field of emotion classification, we have done many experiments under different imbalanced ratios. The results demonstrate that the suggested weighting technique can enhance the performance of SVM even more.
@@
The remaining sections of this paper are structured in the following manner. Section 2 provides a concise overview of the recent literature on the classification of imbalanced data. Section 3 presents the introduction of a novel adaptive cost-sensitive learning and recursive denoising technique. Section 4 involves conducting experiments on various datasets by integrating different algorithms with our suggested technique. The discussion and conclusions can be found in section 5.
@@ 2.1 Imbalanced Data
Data imbalances occur in the real world, often leading to overfitting. Faced with the problem of imbalance in SVM, there are two popular solutions. The first is to reconstruct the training points. Increase the number of points of a few classes in different ways (oversampling) or decrease the number of points of a majority class (undersampling) [20, 21]. The second is to introduce appropriate weight coefficients into the loss function. The error cost or the decision threshold of imbalanced data is adjusted to increase the weight of the minority class and reduce the weight of the majority class to achieve balance. An initial approach to incorporating weighting into SVM is through the use of fuzzy support vector machine (FSVM) [22]. Fuzzy SVM assigns different fuzzy membership degrees to all samples in the training set, and retrains the classifier. Shao et al.[23] introduced an extremely efficient weighted Lagrangian twin support vector machine (WLTSVM) that utilizes distinct training points to create two nearby hyperplanes for the classification of unbalanced data. Fan et al. [24] proposed an improved method based on FSVM to determine fuzzy membership by class determinism of samples, so that distinct samples can provide different contributions to the classification hyperplane.Kang et al. [25] introduced a downsampling method for Support Vector Machines (SVM) that uses spatial geometric distances. This method assigns weights to the samples depending on their Euclidean distance from the hyperplane. The weights highlight the samples’ influence on the decision hyperplane. In order to improve the performance of SVM to generate samples, paper [26] combines oversampling with particle swarm optimization algorithm to find the best sample position.
@@ 3.1 Support Vector Machine(SVM)
$\mathop{\min}_{\mathbf{w}\in\mathbb{R}^{n}}F(\mathbf{w})=\cfrac{1}{l}\sum\limits_{i=1}^{l}f(\mathbf{w},\theta)=\cfrac{C}{2}\lVert\mathbf{w}\rVert^{2}+\cfrac{1}{l}\sum\limits_{i=1}^{l}l((\phi(x_{i}),y_{i});\mathbf{w})$ (3)
where $f(\mathbf{w},\theta_{i})=\cfrac{C}{2}\lVert\mathbf{w}\rVert^{2}+\text{max}(0,1-y_{i}(\mathbf{w^{T}}\phi(x_{i})+b))$, and $C>0$ is a hyperparameter.
@@ 3.2 Soft-Margin Weighted Support Vector Machine (WSVM)
$\mathop{\min}_{\mathbf{w}\in\mathbb{R}^{n}}F(\mathbf{w})=\dfrac{1}{l}\sum\limits_{i=1}^{l}f(\mathbf{w},\theta)=\dfrac{C}{2}\lVert\mathbf{w}\rVert^{2}+\dfrac{1}{l}\sum\limits_{i=1}^{l}l((\phi(x_{i}),y_{i},\alpha_{i});\mathbf{w})$ (4)
where $f(\mathbf{w},\theta_{i})=\frac{1}{2}C\lVert\mathbf{w}\rVert^{2}+\text{max}(0,\alpha_{i}(1-y_{i}(\mathbf{w^{T}}\phi(x_{i})+b)))$
We no longer focus on the overall loss, but on the loss of each sample. In other words, we can regard the weight as a penalty term. The greater the punishment for the sample, the less the corresponding loss is required. In fact, the SVM is more accurate and targeted by punishing each sample separately by setting the weight. It satisfies the spacing maximization principle.
In order to make the expression of WSVM more concise, we give the following explanation
A $:=[\alpha_{1},\alpha_{2},...,\alpha_{l}]^{T}\in\mathbb{R}^{l}$ (5)
1 $:=[1,1,...,1]^{T}\in\mathbb{R}^{l}$
y $:=[y_{1},y_{2},...,y_{l}]^{T}\in\mathbb{R}^{l}$
Z $:=[y_{1}\phi(x_{1}),y_{2}\phi(x_{2}),...,y_{l}\phi(x_{l})]^{T}\in\mathbb{R}^{l\times n}$
Therefore, problem (4) can be expressed more concisely as follows
$\mathop{\min}_{\mathbf{w}\in\mathbb{R}^{n}}F(\mathbf{w})$ $=\frac{C}{2}\lVert\mathbf{w}\rVert^{2}+\frac{1}{l}\textbf{A}^{T}(\textbf{1}-\textbf{Z}\mathbf{w}-b\textbf{y})$ (6)
@@ 3.3 Adaptive Weight Function(AW Function)
The initial decision hyperplane $\mathbf{w_{0}}\phi(x)+b=0$ is obtained. Set $d_{i}^{k}$ is the distance between the $i$ th sample and the initial decision hyperplane, in the $k$ th iteration. In order to summarize the expression of $\mathbf{d_{i}^{k}}$, we give several new definitions.
We can easily get different kinds of clustering points in the kernel space. Set the total number of positive sample population as $\mathbf{n_{+}}$ and its clustering point as $\boldsymbol{\xi_{+}}$. In addition, the total number of negative sample groups is recorded as $\mathbf{n_{-}}$, and the clustering point is recorded as $\boldsymbol{\xi_{-}}$. The expressions of $\boldsymbol{\xi_{+}}$ and $\boldsymbol{\xi_{-}}$ are as follows:
$\boldsymbol{\xi_{+}}=\frac{1}{n_{+}}\sum\limits_{s=1}^{n_{+}}\phi(x_{s})$ (8)
$\boldsymbol{\xi_{-}}=\frac{1}{n_{-}}\sum\limits_{t=1}^{n_{-}}\phi(x_{t})$ (9)
Then the normal vector $\mathbf{w}$ of the decision hyperplane can be expressed as
$\mathbf{w}=\boldsymbol{\xi_{+}}-\boldsymbol{\xi_{-}}$ (10)
In order to calculate the distance between the sample in the kernel space and the decision hyperplane, we set a point on the hyperplane, which is marked as $\boldsymbol{\xi}=\frac{1}{2}(\boldsymbol{\xi_{+}}+\boldsymbol{\xi_{-}})$. Finally, we give the expression of $\mathbf{d_{i}^{k}}$.
$\mathbf{d_{i}^{k}}$ $=\frac{|(\phi(x_{i})-\boldsymbol{\xi})\mathbf{w_{k}}|}{\lVert\mathbf{w_{k}}\rVert}$ (11)
$=\frac{|\mathbf{w_{k}}^{T}(\phi(x_{i})-\boldsymbol{\xi})|}{\lVert\mathbf{w_{k}}\rVert}=\frac{|\mathbf{w_{k}}^{T}\phi(x_{i})-\mathbf{w_{k}}^{T}\boldsymbol{\xi}|}{\lVert\mathbf{w_{k}}\rVert}$
$=\frac{\big{|}\big{(}\boldsymbol{\xi_{+}}-\boldsymbol{\xi_{-}}\big{)}^{T}\phi(x_{i})-\frac{1}{2}\big{(}\boldsymbol{\xi_{+}}-\boldsymbol{\xi_{-}}\big{)}^{T}(\boldsymbol{\xi_{+}}+\boldsymbol{\xi_{-}})\big{|}}{\lVert\mathbf{w_{k}}\rVert}$
$=\frac{\Bigg{|}\frac{1}{n_{+}}\sum\limits_{s=1}^{n_{+}}k(x_{s},x_{i})-\frac{1}{n_{-}}\sum\limits_{t=1}^{n_{-}}k(x_{t},x_{i})}{\lVert\mathbf{w_{k}}\rVert}-\frac{\frac{1}{2}\Big{[}\frac{1}{n_{+}^{2}}\sum\limits_{s_{1}=1}^{n_{+}}\sum\limits_{s_{2}=1}^{n_{+}}k(x_{s_{1}},x_{s_{2}})-\frac{1}{n_{-}^{2}}\sum\limits_{t_{1}=1}^{n_{-}}\sum\limits_{t_{2}=1}^{n_{-}}k(x_{t_{1}},x_{t_{2}})\Big{]}\Bigg{|}}{\lVert\mathbf{w_{k}}\rVert}$
@@
$\boldsymbol{\alpha_{ki}}=\frac{2}{\sqrt{2\pi}\sigma}e^{-\frac{(d_{i}^{k})^{2}}{2\sigma^{2}}}+\frac{1}{M-m}e^{-\frac{d_{i}^{k}}{M-m}}$ (12)
@@
Figure 4: AW function with different $\sigma$ ($\sigma\in\{0.5,1,1.5\}$)
@@ 3.4 Generalized Framework for AW-WSVM
All samples are initially assigned the same weight. The assumption in Section 3.3 is given to illustrate the rationality of the above statement.
Now, we can utilize the AW function mentioned before to produce weights for training the AW-WSVM algorithm. Enter the distance $\mathbf{d_{i}^{k}}(i=1,2,...,l)$ obtained by the $kth$ training into the (12). We filter the data by recursively finding the nearest neighbor. Then update the weight of each sample $\alpha_{ki},i=1,...,l$. Finally, we put it in the objective function again. The framework includes the following steps:
@@
(2) Filter samples
The minority is more sensitive to noise. We can calculate the distance between each sample and the hyperplane, and only pay attention to the samples near the decision boundary (set the threshold according to experience). Because the samples far away from the decision boundary will not have a great influence on the optimal hyperplane. Remember, the majority sample set around the decision hyperplane is $S_{+}$, and the minority sample set is $S_{-}$. Traverse the samples in $S_{-}$ and find the nearest neighbor of $x_{i}$. If the nearest neighbor belongs to $S_{+}$, it means $x_{j}$ is the noise to be removed. We describe this process in detail in Fig. 5;
Figure 5: The process of filtering samples
@@
1:The training set $(x_{i},y_{i},\alpha_{ki}),i=1,…,l$, the initial weight $\alpha_{0}$, the original $\mathbf{w_{0}}=[0,…,0]^{T}$ and $k_{max}$.
@@
6:     Calculate $d_{i}^{k}$ by (11)
7:     Set $S_{+}$ and $S_{-}$
8:     for each $x_{i}\in S_{-}$ do
9:         Calculate the nearest neighbor of $x_{i}$ and express it as $x_{j}$
10:         if $x_{j}\in S_{+}$ then
11:              Eliminate noise
12:         end if
13:     end for
14:     Set
@@
15:     Update $\alpha_{k}=[\alpha_{k1},…,\alpha_{kl}]^{T}$ using (12)
16:     $k\leftarrow k+1$
17:end while
We can draw a conclusion that the normal vector of the decision hyperplane in each iteration can be solved by most stochastic optimization algorithms such as Algorithm 1, 2 or 3. In the case of Algorithm 3 oNAQ, we can easily provide important steps:
$\nabla F_{1}\leftarrow\nabla F(\mathbf{w_{k}}+\mu v_{k},X_{k})=C(\mathbf{w_{k}}+\mu v_{k})-\frac{1}{l}Z^{T}A_{k}$ (13)
$\nabla F_{2}\leftarrow\nabla F(\mathbf{w_{k+1}},X_{k})=C\mathbf{w_{k+1}}-\frac{1}{l}Z^{T}A_{k+1}$ (14)
(a) Weights of Different Samples
(b) 3D Visualisation of Weights
Figure 6: Weights generated by AW function